OpenStudy (mr._scampy):

How would I find [(x^2 - 3) / (x^2 + 2x - 8)] - [(x - 4) / (x + 4)]? I have the answer, but I'd really like to see the steps.

OpenStudy (anonymous):

[(x^2-3)/(x^2+2x-8)] - [(x-4)/(x+4)]=? [(x^2-3)/(x+4)(x-2)] - [(x-4)/(x+4)]=? [(x^2-3)/(x-2)] - [x-4]=? [(x^2-3)/(x-2)] - [x-4]=?

OpenStudy (mr._scampy):

Perfect, thank you